# Can LLMs Reason Structurally? An Evaluation via the Lens of Data Structures

### v2

@@ 4.3 Can LLMs reason structurally on natural language tasks?
While the previous sections evaluated LLMs on canonical data structures, real-world use cases are often expressed in natural language. To bridge this gap, we extend the benchmark with the natural suite, which embeds data structure tasks in narrative, real-world contexts, allowing us to test whether LLMs generalize structural reasoning beyond formal descriptions.
@@
Example natural language prompt for queue compound. On a sunny afternoon in the park, an ice cream truck rolled in… Children began to form a line, each newcomer taking their place at the end while the vendor served from the front… Isabella Miller ran over and joined the ice cream line. The next kid in line was served promptly. Q: What is the order of the remaining kids in line?
Table 7: Model performance on formal and natural descriptors.
Model Task Formal Natural
GPT-4.1 Queue 1.00 0.77
BST 0.88 0.59
Graph 0.42 0.43
o4-mini Queue 1.00 0.83
BST 1.00 0.93
Graph 0.84 0.67
@@
We design three real-world scenarios that implicitly require data structures: queue (children buying ice cream), binary search tree (clinic appointments), and graph (galaxy traveling game). Synthetic data follows the same distributions as Section˜3.3, with realistic substitutions (e.g., names for integers). Each scenario was written by humans and paraphrased by GPT-4o. All prompts were reviewed by three annotators for clarity and unambiguity. Details are supplemented in Section˜A.8.
@@
LLMs struggle when shifting from formal to natural language.
@@
As shown in Table˜7, performance drops when tasks are described in natural language compared to formal descriptions, despite identical problem distribution. The higher accuracy on formal descriptors may stem from training on textbook-style patterns, where integers and explicit syntax are common. This observation suggests that even reasoning models struggle to apply reasoning in language-rich, real-world contexts. Bridging this gap is crucial for reliable deployment and presents a key direction for future research.
4.4 Can LLMs reason structurally with code generation?
@@
As motivated in Section˜1 and Section˜2, our benchmark targets general reasoning independent of code execution or tool use. Nonetheless, to assess whether code generation provides any benefit, we run ablations on six models using the code suite across three modes: (i) CodeOnly, where models generate Python code executed by an external interpreter; (ii) CodeEnforce, where models must write code and reason through its execution internally without relying on an interpreter; and (iii) CodeMaybe, similar to CodeEnforce but makes code generation optional. Full details on the experimental setup and results are provided in Section˜A.9.
@@
Table 9: Individual scores on three data structures under CodeOnly.
o4- mini Gemini- 2.5-pro Claude-3. 7-Sonnet
Geom Graph 0.99 0.98 0.96
DAWG 0.56 0.90 0.89
Graph-Natural 0.69 0.56 0.23
@@
In Table˜9, performance in CodeMaybe and CodeEnforce matches the default setup, showing that writing code offers little benefit over natural language reasoning when models must internally simulate execution. This reinforces our central claim: LLMs remain limited in their ability to perform structural reasoning, even when guided by their own code.
@@
As shown in Table˜9, with an external interpreter in CodeOnly, models perform well on Geom Graph, a standard structure in computer graphics with widely available implementations. In contrast, they struggle on the less familiar DAWG, where custom constraints enforce unambiguous outputs, suggesting reliance on memorized solutions rather than genuine reasoning. Performance drops further on Graph-Natural, where models default to brittle pattern matching (e.g., rigidly mapping “A space tunnel links A and B” to “G.add_edge(A, B)”) but failing to understand paraphrases like “Couriers frequently travel the tunnel connecting A to B.” These results highlight the fragility of structural reasoning under natural language ambiguity, even with code generation and external execution.
@@ 5 Discussion and conclusion
Can LLMs reason structurally? Through DSR-Bench, we provide a systematic answer: not yet. Instruction-tuned models struggle with multi-attribute reasoning (e.g., database indexing) and multi-hop reasoning (e.g., trip planning), while reasoning models achieve only 0.498 accuracy on complex structures and can ignore user-defined constraints. These limitations highlight the need for architectures that support precise function computation, memory mechanisms, and the flexibility to adapt to personalized requests. Evaluations on high-dimensional data (spatial) and natural language scenarios (natural) further reveal gaps between current reasoning capabilities and real-world readiness. Code generation modes (code) show that models cannot reliably reason over their own code, often reverting to memorized patterns or brittle mappings even with external execution.
DSR-Bench provides a systematic framework for evaluating algorithmic reasoning through the lens of structural reasoning. It provides the community with a powerful diagnostic tool: researchers can pinpoint failure modes, test targeted improvements, and measure progress on specific relationship types, paving the way for algorithm-centric model design as an alternative to scaling alone. It also raises new questions: Can LLMs dynamically choose reasoning strategies? Where and why do errors arise in intermediate steps? We invite the community to use DSR-Bench to explore these directions.
Acknowledgments
@@
Appendix A Appendix
A.1 Additional related work on LLM benchmarking
@@
Large language models (LLMs) have demonstrated remarkable performance across a wide range of applications, prompting growing interest in understanding their capabilities and limitations. Recent efforts have focused on systematically benchmarking LLMs on core natural language processing tasks, including language understanding (Wang et al., 2024; Feng et al., 2024), text generation (Maynez et al., 2023; Singh et al., 2024), reasoning (Feng et al., 2024; Valmeekam et al., 2023), and machine translation (Yao et al., 2023). Additionally, some benchmarks evaluate ethical dimensions such as robustness, bias, and trustworthiness (Siska et al., 2024; Luo et al., 2024; Koo et al., 2023; Li et al., 2024b; Zhang et al., 2024b; Liu et al., 2023c). Specialized benchmarks have also emerged in scientific and technical domains, including mathematics (White et al., 2025; Liu et al., 2024a), programming (White et al., 2025; Jain et al., 2025; Chen et al., 2021; Zheng et al., 2023), data analysis (Sui et al., 2024; Liu et al., 2023b; White et al., 2025; Tang et al., 2023; Jiang et al., 2023), and medicine (Li et al., 2024a; Liu et al., 2023a; Longwell et al., 2024; Xu et al., 2024b; Kanithi et al., 2024).
While several studies examine how LLMs convert unstructured inputs into tabular or relational formats (Tang et al., 2023; Jiang et al., 2023), our work explores a distinct question: How well can LLMs construct, manipulate, and reason about classic data structures such as stacks, trees, and graphs? To our knowledge, this is the first comprehensive benchmark targeting this capability, offering a conceptually different evaluation from prior work on structured data.
A.2 Details of data structures and operations
@@
In this section, we list the data structures and the corresponding operations tested in Table˜10. We then provide detailed descriptions of each data structure and explain how we specify their implementations to eliminate ambiguity.
@@
Table 10: Summary of data structures and associated operations in DSR-Bench. Data structures marked with * are included in the challenge suite. All compound operations without explicit specification consist of (insert, delete).
@@
A.3 Examples of prompting strategies
@@
A.4 Accuracy by task and length level across all models
In this section, we provide supplementary accuracy tables for all models in DSR-Bench, broken down by task and length level. Table˜11 summarizes the accuracy of instruction-tuned models on a subset of basic data structures across different length levels. Table˜12 presents the accuracy of reasoning models on selected data structures from the DSR-Bench-challenge suite. For detailed per-model results across all tasks and length levels, see  Table˜13 (GPT-5),  Table˜14 (o4-mini),  Table˜15 (Gemini-2.5-Pro),  Table˜16 (Claude-3.7-Sonnet), Table˜17 (DeepSeek-R1),  Table˜18 (GPT-4.1),  Table˜19 (Gemini-2.0-Flash),  Table˜20 (Claude-3.5-Sonnet),  Table˜21 (DeepSeek-V3), and  Table˜22 (Llama-3.3).
@@ A.4 Accuracy by task and length level across all models
Table 12: Average accuracy on the DSR-Bench-challenge suite for reasoning models (3 runs, scaled to [0, 1], rounded to two decimals).
Category Data Structure Length o4-mini Gemini-2.5 Pro Claude-3.7 Sonnet Deepseek R1 GPT-5 (med)
Temporal Priority Queue Short 0.89 0.89 0.70 0.92 0.84
Medium 0.47 0.41 0.11 0.54 0.44
Long 0.30 0.23 0.04 0.48 0.28
Associative Trie Short 0.99 0.93 0.23 0.93 0.97
Medium 0.73 0.56 0.00 0.50 0.94
Long 0.32 0.37 0.00 0.05 0.92
Suffix Tree Short 0.96 0.96 0.96 1.00 1.00
Medium 0.87 0.89 0.98 0.98 1.00
Long 0.37 0.86 0.79 0.90 0.95
Skip List Short 0.84 0.94 0.87 0.89 0.88
Medium 0.60 0.87 0.76 0.63 0.66
Long 0.41 0.53 0.61 0.54 0.51
Hierarchical Heap Short 0.61 0.57 0.70 0.63 0.78
Medium 0.72 0.28 0.37 0.53 0.71
Long 0.71 0.29 0.13 0.27 0.71
Red Black Tree Short 0.92 0.91 0.69 0.86 0.92
Medium 0.67 0.47 0.11 0.63 0.79
Long 0.37 0.08 0.12 0.37 0.59
B+ Tree Short 0.99 1.00 0.70 0.49 0.97
Medium 0.98 0.97 0.32 0.23 0.99
Long 0.94 0.94 0.13 0.21 0.98
K-D Heap Short 0.22 0.23 0.11 0.23 0.23
Medium 0.09 0.06 0.03 0.08 0.08
Long 0.00 0.00 0.00 0.01 0.00
K-D Tree Short 0.59 0.96 0.00 1.00 0.97
Medium 0.43 0.64 0.00 0.34 0.90
Long 0.38 0.16 0.00 0.01 0.67
Network DSU Short 1.00 1.00 1.00 1.00 1.00
Medium 1.00 0.99 1.00 1.00 1.00
Long 0.94 0.90 0.98 0.99 1.00
Hybrid Bloom Filter Short 1.00 0.94 0.44 0.99 1.00
Medium 0.77 0.97 0.03 0.92 0.84
Long 0.07 0.66 0.00 0.31 0.47
DAWG Short 0.49 0.61 0.17 0.40 1.00
Medium 0.20 0.08 0.00 0.02 0.03
Long 0.06 0.01 0.00 0.00 0.00
Geom Graph Short 0.37 0.19 0.07 0.36 0.21
Medium 0.10 0.17 0.00 0.01 0.08
Long 0.00 0.02 0.00 0.01 0.19
A.4.1 Performance of GPT-5
Table 13: Mean (± std) accuracy of GPT-5 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@ A.4.1 Performance of GPT-5
Associative Hashmap Compound 1.00 (0.00) 0.89 (0.02) 0.71 (0.10)
Trie* Compound 0.97 (0.03) 0.94 (0.05) 0.92 (0.07)
Suffix Tree* Construct 1.00 (0.00) 1.00 (0.00) 0.95 (0.02)
@@
A.4.2 Performance of o4-mini
Table 14: Mean (± std) accuracy of o4-mini on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@ A.4.2 Performance of o4-mini
Associative Hashmap Compound 0.89 (0.04) 0.37 (0.00) 0.26 (0.08)
Trie* Compound 0.99 (0.02) 0.73 (0.06) 0.32 (0.05)
Suffix Tree* Construct 0.96 (0.02) 0.87 (0.03) 0.37 (0.07)
@@
A.4.3 Performance of Gemini-2.5-Pro
Table 15: Mean (± std) accuracy of Gemini-2.5-Pro on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@
Associative Hashmap Compound 0.58 (0.02) 0.16 (0.07) 0.11 (0.04)
Trie* Compound 0.93 (0.03) 0.56 (0.02) 0.37 (0.03)
Suffix Tree* Construct 0.96 (0.04) 0.89 (0.02) 0.86 (0.08)
@@
A.4.4 Performance of Claude-3.7-Sonnet
Table 16: Mean (± std) accuracy of Claude-3.7-Sonnet on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.
@@
Associative Hashmap Compound 0.71 (0.02) 0.16 (0.05) 0.04 (0.05)
Trie* Compound 0.94 (0.02) 0.64 (0.07) 0.31 (0.10)
Suffix Tree* Construct 0.23 (0.00) 0.00 (0.00) 0.00 (0.00)
@@
A.4.5 Performance of DeepSeek-R1
Table 17: Mean (± std) accuracy of DeepSeek-R1 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge suite.

### v3

@@ 4.3 Can LLMs reason structurally on realistic tasks?
While the previous sections evaluated LLMs on canonical data structures, real-world use cases are often described in messy, context-rich language. To bridge this gap, we extend the benchmark with the realistic probe, which embeds data structure tasks in narrative, real-world contexts, allowing us to test whether LLMs generalize structural reasoning beyond formal descriptions.
@@
Example prompt for queue in the realistic probe. On a sunny afternoon in the park, an ice cream truck rolled in… Children began to form a line, each newcomer taking their place at the end while the vendor served from the front… Isabella Miller ran over and joined the ice cream line. The next kid in line was served promptly. Q: What is the order of the remaining kids in line?
@@
We design three real-world scenarios that implicitly require data structures: queue (children buying ice cream), binary search tree (clinic appointments), and graph (galaxy traveling game). Synthetic data follows the same distributions as Section˜3.4, with realistic substitutions (e.g., names for integers). Each scenario was written by humans and paraphrased by GPT-4o. All prompts were reviewed by three annotators for clarity and unambiguity. Details are provided in Appendix˜G.
Figure 6: Model performance on formal and realistic usages.
@@
LLMs struggle when shifting from formal to realistic usage.
@@
Performance drops when tasks are using realistic, context-rich language compared to formal descriptors, despite identical problem distribution (Figure˜6). The higher accuracy on formal descriptors may stem from training on textbook-style patterns, where integers and explicit syntax are common. This observation suggests that even reasoning models struggle to apply reasoning in realistic contexts. Bridging this gap is crucial for reliable deployment and presents a key direction for future research.
4.4 Can LLMs reason structurally with code?
@@
As motivated in Sections˜1 and 2, our benchmark targets LLMs’ inherent reasoning independent of code execution or tool use. Nonetheless, to assess whether code generation provides any benefit, we run ablations on six models using the code probe across three modes: (i) CodeOnly, where models generate Python code executed by an external interpreter; (ii) CodeEnforce, where models must write code and reason through its execution internally without relying on an interpreter; and (iii) CodeMaybe, similar to CodeEnforce but makes code generation optional. Full details on the experimental setup and results are provided in Appendix˜H.
@@
In Table˜4, performance in CodeMaybe and CodeEnforce matches the default setup, despite writing high-quality code as shown in CodeOnly. This observation suggests that writing code offers little benefit over natural language reasoning when models must internally simulate execution. This reinforces our claim: LLMs remain limited in their ability to perform structural reasoning, even when guided by their own code.
Figure 7: Scores for CodeOnly.
@@
As shown in Figure˜7, with an external interpreter in CodeOnly, models perform well on Geom Graph, a standard structure in computer graphics with widely available implementations online. In contrast, they struggle on the less familiar DAWG, where custom constraints enforce unambiguous outputs, suggesting reliance on memorized solutions rather than genuine reasoning. Performance drops further on Graph-Natural, where models default to brittle pattern matching (e.g., rigidly mapping “A tunnel links A and B” to “G.add_edge(A, B)”) but failing to understand paraphrases like “Couriers frequently travel the tunnel connecting A to B.” These results highlight the fragility of structural reasoning in the presence of context-rich language, even with code generation and external execution.
@@ 5 Discussion and conclusion
Can LLMs reason structurally? Through DSR-Bench, we provide a systematic answer: not yet. Instruction-tuned models struggle with multi-attribute reasoning (e.g., database indexing) and multi-hop reasoning (e.g., trip planning), while reasoning models achieve only 0.46 accuracy on complex structures and can ignore user-defined constraints. These limitations highlight the need for architectures that support precise function computation, memory mechanisms, and the flexibility to adapt to user-defined constraints. Evaluations on high-dimensional data (spatial) and context-rich scenarios (realistic) further reveal gaps between current reasoning capabilities and real-world deployment. Code generation modes (code) show that models cannot reliably reason about their own code, often reverting to memorized patterns or brittle mappings even with external execution.
Outlook
DSR-Bench provides a systematic framework for evaluating algorithmic reasoning through the lens of data structures. It provides the community with a powerful diagnostic tool: researchers can pinpoint failure modes, test targeted improvements, and measure progress on specific relationship types, paving the way for algorithm-centric model design (Eberle et al., 2025). It also raises new questions: Can LLMs dynamically choose reasoning strategies? Can multimodal LLMs reason over DSR-Bench tasks using visual inputs? Furthermore, it is well-suited as a testbed for mechanistic interpretability and RL fine-tuning due to its synthetic and verifiable nature. We invite the community to use DSR-Bench to explore these future directions.
Acknowledgements
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Additional related works
LLM Benchmarking
@@ Appendix A Additional related works
LLMs have demonstrated remarkable performance across a wide range of applications, prompting growing interest in understanding their capabilities and limitations. Recent efforts have focused on systematically benchmarking and evaluating LLMs on core natural language processing tasks, including language understanding (Wang et al., 2024b), text generation (Maynez et al., 2023; Singh et al., 2024), reasoning (Valmeekam et al., 2023; Wang et al., 2024a), and machine translation (Yao et al., 2024). Additionally, some benchmarks evaluate alignment and safety, such as robustness, bias, and trustworthiness (Siska et al., 2024; Luo et al., 2025; Koo et al., 2024; Li et al., 2024b; Zhang et al., 2024b; Liu et al., 2024c). Specialized benchmarks have also emerged in scientific and technical domains, including mathematics (White et al., 2025; Liu et al., 2024a), programming (White et al., 2025; Jain et al., 2025; Chen et al., 2021; Zheng et al., 2023), data analysis (Sui et al., 2024; Liu et al., 2023b; White et al., 2025; Tang et al., 2024; Jiang et al., 2023), and medicine (Li et al., 2024a; Liu et al., 2023a; Xu et al., 2024b; Kanithi et al., 2026). While several studies examine how LLMs convert unstructured inputs into tabular or relational formats (Tang et al., 2024; Jiang et al., 2023), our work explores a distinct question: How well can LLMs construct, manipulate, and reason about classic data structures such as stacks, trees, and graphs? To our knowledge, this is the first comprehensive benchmark targeting this capability, offering a conceptually different evaluation from prior work on structured data.
Reasoning benchmarks
Existing LLM reasoning benchmarks are predominantly high-level and domain-specific, targeting math (Cobbe et al., 2021; Liu et al., 2024a, b), STEM (Hendrycks et al., 2021), and logic puzzles (White et al., 2025; Giadikiaroglou et al., 2024). These often require complex responses with intertwined reasoning steps, relying on subjective human or LLM-based evaluation (Chiang et al., 2024; Feuer et al., 2025; Ye et al., 2025). We focus on structural reasoning, an implicit requirement underlying problem-solving across many domains, and use it as a framework to evaluate algorithmic reasoning. By using data structures as clear abstractions of different data relationships, we isolate algorithmic reasoning from domain-specific complexities.
Coding benchmarks
Coding benchmarks evaluate how well LLMs write syntactically correct code or function as coding agents (Chen et al., 2021; Zheng et al., 2023; Jimenez et al., 2024; Jain et al., 2025; White et al., 2025; Aider-AI, 2025), typically requiring external interpreters for verification. While useful, these benchmarks conflate reasoning with tool execution and are limited to domains where coding applies. Coding tasks for algorithmic reasoning are also susceptible to data contamination due to the abundance of online coding materials for algorithmic tasks. In contrast, we target LLMs’ inherent reasoning skills independent of external tools, reflecting the broader goal of assessing progress toward general intelligence. Prior works (Malfa et al., 2024, 2025; Liu et al., 2025) study code simulation as a lens to probe general reasoning. Instead, we specifically focus on structural reasoning via data structure tasks and include code to probe whether code generation aids this process.
Appendix B Details of data structures and operations
@@
In this section, we list the data structures and the corresponding operations tested in Table˜5. We then provide detailed descriptions of each data structure and explain how we specify their implementations to eliminate ambiguity.
@@
Table 5: Summary of data structures and associated operations in DSR-Bench. Data structures marked with * are included in the challenge subset. All compound operations without explicit specification consist of (insert, delete).
@@
Appendix C Examples of prompting strategies
@@
Appendix D Accuracy by task and length level across all models
In this section, we provide supplementary accuracy tables for all models in DSR-Bench, broken down by task and length level. Table˜6 summarizes the accuracy of instruction-tuned models on a subset of basic data structures across different length levels. For detailed per-model results across all tasks and length levels, see  Table˜7 (GPT-5),  Table˜8 (o4-mini),  Table˜9 (Gemini-2.5-Pro),  Table˜10 (Claude-3.7-Sonnet), Table˜11 (DeepSeek-R1),  Table˜12 (GPT-4.1),  Table˜13 (Gemini-2.0-Flash),  Table˜14 (Claude-3.5-Sonnet),  Table˜15 (DeepSeek-V3), and  Table˜16 (Llama-3.3).
@@ Appendix D Accuracy by task and length level across all models
D.1 Performance of GPT-5
Table 7: Mean (± std) accuracy of GPT-5 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@ D.1 Performance of GPT-5
Associative Hashmap* Compound 1.00 (0.00) 0.89 (0.02) 0.71 (0.10)
Trie Compound 0.97 (0.03) 0.94 (0.05) 0.92 (0.07)
Suffix Tree Construct 1.00 (0.00) 1.00 (0.00) 0.95 (0.02)
@@
D.2 Performance of o4-mini
Table 8: Mean (± std) accuracy of o4-mini on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@ D.2 Performance of o4-mini
Associative Hashmap* Compound 0.89 (0.04) 0.37 (0.00) 0.26 (0.08)
Trie Compound 0.99 (0.02) 0.73 (0.06) 0.32 (0.05)
Suffix Tree Construct 0.96 (0.02) 0.87 (0.03) 0.37 (0.07)
@@
D.3 Performance of Gemini-2.5-Pro
Table 9: Mean (± std) accuracy of Gemini-2.5-Pro on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@
Associative Hashmap* Compound 0.58 (0.02) 0.16 (0.07) 0.11 (0.04)
Trie Compound 0.93 (0.03) 0.56 (0.02) 0.37 (0.03)
Suffix Tree Construct 0.96 (0.04) 0.89 (0.02) 0.86 (0.08)
@@
D.4 Performance of Claude-3.7-Sonnet
Table 10: Mean (± std) accuracy of Claude-3.7-Sonnet on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.
@@
Associative Hashmap* Compound 0.71 (0.02) 0.16 (0.05) 0.04 (0.05)
Trie Compound 0.94 (0.02) 0.64 (0.07) 0.31 (0.10)
Suffix Tree Construct 0.23 (0.00) 0.00 (0.00) 0.00 (0.00)
@@
D.5 Performance of DeepSeek-R1
Table 11: Mean (± std) accuracy of DeepSeek-R1 on all DSR-Bench tasks over three runs. Data structures marked with * are included in the DSR-Bench-challenge subset.